# Minimal criteria for continuous-variable genuine multipartite entanglement

### v1

@@ V.1.2 Numerical examples
where the elements were rounded for brevity to one decimal place in which case physicality of the rounded CM remained preserved. Minimization of the differences $D_{P}$ and $D_{S}$ over the parameters $\ell^{\alpha}_{ij}\in[-1,1]$ yields $D_{P}=-0.245$ and $D_{S}=-0.165$, where the respective rounded optimal values of the parameters $\ell^{\alpha}_{ij}$ can be found in Appendix B. As both the differences are negative the GME of the state with CM (61) is detected both by the product criterion (V.1), as well as sum criterion (V.1).
@@
Indeed, while the product criterion (V.1) yields a non-negative minimal difference $D_{P}$ and thus it does not detect GME, for the sum criterion (V.1) one gets $D_{S}=-0.070$, which clearly confirms the presence of GME in the considered state. The rounded optimal values of the parameters $\ell^{\alpha}_{ij}$ attaining the latter difference are given explicitly in Appendix B.
V.1.3 Example derived from a guessed witness
There is yet another method of how one can find examples of three-mode Gaussian states violating inequality (V.1). What is more, below we use the method to find examples of GME states detected by the sum criteria for more than three modes.
Typically, the inseparability criteria based on uncertainty relations are tailored to the state whose entanglement is to be detect, which is know. Here, on the other hand, we know the criteria and seek states, which are detected by them. This task can be solved relatively easily if we set the parameters $\ell^{\alpha}_{ij}$ to suitable values. Namely, by using Eq. (20) and dividing the LHS of the criterion with the RHS, we can express the inequality (50) in the form $\mbox{Tr}[\gamma Z]\geq 1$, where $Z=Z^{x}\oplus Z^{p}$ and $\gamma=\gamma^{x}\oplus\gamma^{p}$. Here, $Z$ is a real $2N\times 2N$ symmetric positive-semidefinite matrix and provided that the parameters $\ell^{\alpha}_{ij}$ are chosen properly, it can detect GME. In fact, the matrix $Z$ is nothing but the already mentioned GME witness in the space of covariance matrices [26].
@@ V.1.3 Example derived from a guessed witness
$SZS^{T}=\mbox{diag}(\nu_{1},\nu_{2},\ldots,\nu_{N},\nu_{1},\nu_{2},\ldots\nu_{N})\equiv Z_{W}.$
@@
The GME witness $Z$ can be obtained, e.g., by setting some parameters $\ell^{\alpha}_{ij}$ equal to $+1$ or $-1$, often small parameters, like $\ell^{\alpha}_{33}$, equal to $10^{-2}$, and by guessing the remaining free parameters. For example, consider the values of the parameters $\ell^{\alpha}_{ij}$ in Tab. 2.
@@
$Z_{3}=\frac{1}{168}\left[\left(\begin{array}[]{ccc}64&-8&0\\ -8&2&-1\\ 0&-1&1.0001\end{array}\right)\oplus\left(\begin{array}[]{ccc}100&20&0\\ 20&29&5\\ 0&5&1.0001\end{array}\right)\right].$ (70)
Next, we find numerically the matrix $S_{3}$ symplectically diagonalizing $Z_{3}$, calculate the CM $\gamma_{3}=S_{3}^{T}S_{3}$, round it for brevity to two decimal places and add to the rounded matrix $0^{-2}\times\openone1$ to get a physical CM (see Appendix C.1 for details of the derivation). Thus we arrive at the CM
$\gamma_{3}=\left(\begin{array}[]{ccc}1.40&1.15&0.30\\ 1.15&7.64&2.48\\ 0.30&2.48&1.49\\ \end{array}\right)\oplus\left(\begin{array}[]{ccc}0.84&-0.15&0.09\\ -0.15&0.33&-0.50\\ 0.09&-0.50&1.49\\ \end{array}\right).$ (78)
@@
V.2 Multi-mode states
@@
Table 3: All non-isomorphic trees with $N=3,4,5$ and $6$ vertices. The root of the tree is depicted by a red circle. The linear trees are labeled according to the standard labeling, whereas the other trees are covered by the reverse level order labeling, which lead to the GME criteria involving variances of at most two-mode quadrature combinations. The last column contains the difference $D_{S}$, Eq. (53), for the displayed trees and CMs constructed by means of the method of Subsec. V.1.3. The values of $D_{S}$ are calculated for the values of parameters $\ell^{\alpha}_{ij}$ found by a numerical minimization over $\ell^{\alpha}_{ij}\in[-10,10]$. The cases for which we did not found an example are denoted by a cross. Explicit form of all the CMs as well as other details of derivation of $D_{S}$ can be found in Appendix C. See text for details.
$N$ Index Labelled tree Difference $D_{S}$
3 $3$ 123 $-17.87$
4 $4a$ 1 2 3 4 $-20.31$
$4b$ 3412 $-5.36$
[…]
$5b$ 52431 $-3.64$
$5c$ 12543 $-2.30$
[…]
$6c$ 2 5 6 3 1 4 $-1.00$
$6d$ 3 5 6 4 1 2 $-0.91$
$6e$ 5 6 1 2 4 3 $-0.65$
$6f$ 5 6 4 2 3 1 $-0.13$
Since optimization over $\ell^{\alpha}_{ij}\in[-1,1]$ yields typically for $N>3$ small values of $D_{S}$, the values of the difference in Tab. 3 were obtained for optimization over an extended interval $\ell^{\alpha}_{ij}\in[-10,10]$. This is why the value of $D_{S}$ for the three-mode case is roughly hundred-times larger compared to the examples presented in previous section. The Tab. 3 indicates that up to the configurations $3$ and $4a$ the value of $D_{S}$ decreases with increasing $N$, which is a typical behaviour of GME states (cf. Tab. I of Ref. [33]). The exception is likely to be attributed to the fact that in the case $3$ we guessed a little worse witness than for configuration $4a$. All the guessed witnesses, the CMs of the states, and other details of derivation of the results in Tab. 3 are given explicitly in Appendix C.
@@
VI Conclusions
In this paper we derived a set of product and sum criteria for GME of continuous-variable systems. The criteria are based on the second moments of quadrature operators and unlike the overwhelming majority of the known GME criteria contain only the variances of at most two-mode quadrature combinations. Moreover, the presented criteria are also minimal in the sense that they contain only the minimum number of the latter variances. In other words, the criteria require only the knowledge of those elements of the CM of the investigated state, which belong to the minimal set of two-mode marginal CMs. Since our criteria do not take into account the $x-p$ correlations, the number of the elements is $24N-$ 2 and thus it scales only linearly with the number of modes. This distinguishes our criteria from other GME criteria which always utilize a strictly larger number of elements of the CM, which scales quadratically with the number of modes. Since any minimal set of marginal CMs can be represented by a special kind of a graph known as a tree, each criterion is intimately associated to a particular tree. This allowed us to develop a method of construction of all the sum criteria directly from the underlying tree.
Further, we have shown that the product criterion detects GME of the three-mode CV GHZ-like state and the split squeezed state, but neither of the states is detected by the sum criterion. We have also constructed numerical examples of three-mode Gaussian GME states which are simultaneously detected by both the criteria as well as only by the sum criterion. To get examples of GME states detected by our criteria for more than three modes, we developed a method based on symplectic diagonalization of a guessed witness in the space of CMs. Thus we found states detected by the sum criterion for nearly all configurations of up to six modes.
@@ VI Conclusions
Acknowledgements.
O.L. acknowledges support from IGA-PrF-2022-005 and the European Union’s 2020 research and innovation program (CSA - Coordination and support action, H2020-WIDESPREAD-2020-5) under grant agreement No. 951737 (NONGAUSS). L.M. acknowledges the project 8C22002 (CVStar) of MEYS of the Czech Republic, which has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement no. 731473 and 101017733.
@@ Appendix A Fully inseparable biseparable Gaussian state
This section is dedicated to the proof that the three-mode biseparable state (13) with CM (14) is in a certain region of squeezing parameters fully inseparable, i.e., it is entangled across all three bipartite splits 1|23,2|311|23,2|311 | 23 , 2 | 31 and $3|12$. The state possesses a fully symmetric CM of the form
@@
[…]
@@
Appendix B Optimal parameters
@@
For the three-mode CM (61) and the product criterion (V.1) the optimal parameters $\ell^{\alpha}_{ij}$ are summarized in Tab. 4
Table 4: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference $D_{P}=-0.245$ for the product criterion (V.1) and the CM $\gamma_{1}$, Eq. (61).
@@
In the case of CM (61) and the sum criterion (V.1) the optimal parameters $\ell^{\alpha}_{ij}$ are summarized in Tab. 5
Table 5: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference $D_{S}=-0.165$ for the sum criterion (V.1) and the CM $\gamma_{1}$, Eq. (61).
@@
Optimal parameters $\ell^{\alpha}_{ij}$ for the CM (62) and the sum criterion (V.1) are summarized in Tab. 6
Table 6: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference $D_{S}=-0.070$ for the sum criterion (V.1) and the CM $\gamma_{2}$, Eq. (62).
@@
$\ell^{x}_{32}$0.908 $\ell^{p}_{32}$ 0.559
$\ell^{x}_{33}$ $3\times 10^{-5}$ $\ell^{p}_{33}$ $8.7\times 10^{-7}$
Appendix C Derivation of examples from a guessed witness

### v3

@@ 5.1.2 Numerical examples
where the elements were rounded for brevity to one decimal place in which case physicality of the rounded CM remained preserved. Minimization of the differences $D_{P}$ and $D_{S}$ over the parameters $\ell^{\alpha}_{ij}\in[-1,1]$ yields $D_{P}=-0.245$ and $D_{S}=-0.165$, where the respective rounded optimal values of the parameters $\ell^{\alpha}_{ij}$ can be found in Appendix C. As both the differences are negative the GME of the state with CM (64) is detected both by the product criterion (5.1), as well as sum criterion (5.1).
@@
Indeed, while the product criterion (5.1) yields a non-negative minimal difference $D_{P}$ and thus it does not detect GME, for the sum criterion (5.1) one gets $D_{S}=-0.070$, which clearly confirms the presence of GME in the considered state. The rounded optimal values of the parameters $\ell^{\alpha}_{ij}$ attaining the latter difference are given explicitly in Appendix C.
5.1.3 Example derived from a guessed witness
There is yet another method of how one can find examples of three-mode Gaussian states violating inequality (5.1). What is more, below we use the method to find examples of GME states detected by the sum criteria for more than three modes.
Typically, the inseparability criteria based on uncertainty relations are tailored to the state whose entanglement is to be detect, which is known. Here, on the other hand, we know the criteria and seek states, which are detected by them. This task can be solved relatively easily if we set the parameters $\ell^{\alpha}_{ij}$ to suitable values. Namely, by using Eq. (4.1) and dividing the LHS of the criterion with the RHS, we can express the inequality (46) in the form $\mbox{Tr}[\gamma Z]\geq 1$, where $Z=Z^{x}\oplus Z^{p}$ and $\gamma=\gamma^{x}\oplus\gamma^{p}$. Here, $Z$ is a real $2N\times 2N$ symmetric positive-semidefinite matrix and provided that the parameters $\ell^{\alpha}_{ij}$ are chosen properly, it can detect GME. In fact, the matrix $Z$ is nothing but the already mentioned GME witness in the space of covariance matrices [33].
@@ 5.1.3 Example derived from a guessed witness
$SZS^{T}=\mbox{diag}(\nu_{1},\nu_{2},\ldots,\nu_{N},\nu_{1},\nu_{2},\ldots\nu_{N})\equiv Z_{W}.$ (66)
@@
The GME witness $Z$ can be obtained, e.g., by setting some parameters $\ell^{\alpha}_{ij}$ equal to +1+1 or 1-1, often small parameters, like $\ell^{\alpha}_{33}$, equal to $10^{-2}$, and by guessing the remaining free parameters. For example, consider the values of the parameters $\ell^{\alpha}_{ij}$ in Tab. 2.
@@
$Z_{3}=$ $\frac{1}{168}\left[\left(\begin{array}[]{ccc}64&-8&0\\ -8&2&-1\\ 0&-1&1.0001\end{array}\right)\right$. (70)
$\quad\left.\oplus\left(\begin{array}[]{ccc}100&20&0\\ 20&29&5\\ 0&5&1.0001\end{array}\right)\right].$ (74)
Next, we find numerically the matrix $S_{3}$ symplectically diagonalizing $Z_{3}$, calculate the CM $\gamma_{3}=S_{3}^{T}S_{3}$, round it for brevity to two decimal places and add to the rounded matrix $10^{-2}\times\mathbbm{1}$ to get a physical CM (see Appendix D.1 for details of the derivation). Thus we arrive at the CM
$\gamma_{3}=$ $\left(\begin{array}[]{ccc}1.40&1.15&0.30\\ 1.15&7.64&2.48\\ 0.30&2.48&1.49\\ \end{array}\right)$ (78)
$\oplus\left(\begin{array}[]{ccc}0.84&-0.15&0.09\\ -0.15&0.33&-0.50\\ 0.09&-0.50&1.49\\ \end{array}\right).$ (82)
@@
5.1.4 Effect of losses
In the last part of this section, we discuss the effect of imperfections that may occur in the possible experimental preparation of some of the investigated three-mode states on the detection of the given state by the product criterion (5.1). The imperfection we focus on takes the form of the transition of the states through a lossy channel with intensity transmissivity $\eta$. Our analysis starts with the three-mode CV GHZ-like state with CM (8). The best loss robustness for the product criterion (5.1) is found for squeezing parameter $r=0.65$ $(-5.65\;\mathrm{dB})$, where minimal transmissivity is approximately $\eta=0.96$ or $4\%$ loss. Better resilience to losses can be found in the case of the numerically found state with CM (64) where the minimal transmissivity is around $\eta=0.92$ or $8\%$ of losses for the form with elements rounded to one decimal place. The original form of CM (64) is even more robust with minimal transsitivity $\eta=0.89$ or $11\%$ of losses. Quadrature squeezing required to prepare the state with the CM (64) is around $10\;\mathrm{dB}$. From the experimental perspective it is desirable to look for states requiring smaller amounts of squeezing for the preparation. In that case, we can find another numerical example of a state with CM given in the Appendix E that requires only $4.46\;\mathrm{dB}$ of squeezing. This state can be detected using the presented criteria (5.1) starting from a minimal transmissivity of the lossy channel $\eta=0.90$ or $10\%$ of losses.
5.2 Multi-mode states
@@
Table 3: All non-isomorphic trees with $N=3,4,5$ and 66 vertices. The root of the tree is depicted by a red circle. The linear trees are labeled according to the standard labeling, whereas the other trees are covered by the reverse level order labeling, which lead to the GME criteria involving variances of at most two-mode quadrature combinations. The last column contains the difference $D_{S}$, Eq. (49), for the displayed trees and CMs constructed by means of the method of Subsec. 5.1.3. The values of $D_{S}$ are calculated for the values of parameters $\ell^{\alpha}_{ij}$ found by a numerical minimization over $\ell^{\alpha}_{ij}\in[-10,10]$. The cases for which we did not found an example are denoted by a cross. Explicit form of all the CMs as well as other details of derivation of $D_{S}$ can be found in Appendix D. See text for details.
NN Index Labelled tree Difference $D_{S}$
3 33 123 17.87-17.87
4 $4a$ 1 2 3 4 20.31-20.31
$4b$ 3412 5.36-5.36
5 $5a$ 12345 $\times$
$5b$ 52431 3.64-3.64
$5c$ 12543 2.30-2.30
6 $6a$ 1 2 3 4 5 6 $\times$
$6b$ 632154 $\times$
$6c$ 2 5 6 3 1 4 1.00-1.00
$6d$ 3 5 6 4 1 2 0.91-0.91
$6e$ 5 6 1 2 4 3 0.65-0.65
$6f$ 5 6 4 2 3 1 0.13-0.13
Since optimization over $\ell^{\alpha}_{ij}\in[-1,1]$ yields typically for $N>3$ small values of $D_{S}$, the values of the difference in Tab. 3 were obtained for optimization over an extended interval $\ell^{\alpha}_{ij}\in[-10,10]$. This is why the value of $D_{S}$ for the three-mode case is roughly hundred-times larger compared to the examples presented in previous section. The Tab. 3 indicates that up to the configurations $3$ and $4a$ the value of $D_{S}$ decreases with increasing $N$, which is a typical behaviour of GME states (cf. Tab. I of Ref. [41]). The exception is likely to be attributed to the fact that in the case $3$ we guessed a little worse witness than for configuration $4a$. All the guessed witnesses, the CMs of the states, and other details of derivation of the results in Tab. 3 are given explicitly in Appendix D.
@@ 6 $6a$ 1 2 3 4 5 6 $\times$
6 Conclusions
In this paper we derived a set of product and sum criteria for GME of continuous-variable systems. The criteria are based on the second moments of quadrature operators and unlike the overwhelming majority of the known GME criteria contain only the variances of at most two-mode quadrature combinations. Moreover, the presented criteria are also minimal in the sense that they contain only the minimum number of the latter variances. In other words, the criteria require only the knowledge of the minimal set of two-mode marginal CMs, which consists of 7N47N-4 independent CM elements of the state under study [43]. Since our criteria do not take into account the xpx-p correlations, the number of the elements further reduces to 4N24N-2 and thus it scales only linearly with the number of modes. This distinguishes our criteria from the most commonly used GME criteria [38, 36, 37], which require knowledge of the entire CM without xpx-p correlations, i.e., knowledge of its $N(N+1)$ elements. Other known $N$-mode CV GME criteria [33] utilize the whole CM including xpx-p correlations, i.e., to use them one needs to know its $N(2N+1)$ elements. All these criteria thus scale quadratically with the number of modes. Since any minimal set of marginal CMs can be represented by a special kind of a graph known as a tree, each criterion is intimately associated to a particular tree. This allowed us to develop a method of construction of all the sum criteria directly from the underlying tree.
Further, we have shown that the product criterion detects GME of the three-mode CV GHZ-like state and the split squeezed state, but neither of the states is detected by the sum criterion. The fact that we detected GME in the three-mode CV GHZ-like state from its two-mode marginals is in contrast with results for qubit GHZ state where this is not possible [44]. We have also constructed numerical examples of three-mode Gaussian GME states which are simultaneously detected by both the criteria as well as only by the sum criterion. To get examples of GME states detected by our criteria for more than three modes, we developed a method based on symplectic diagonalization of a guessed witness in the space of CMs. Thus we found states detected by the sum criterion for nearly all configurations of up to six modes.
@@ 6 Conclusions
O.L. acknowledges support from IGA-PrF-2023-006 and the European Union’s 2020 research and innovation program (CSA - Coordination and support action, H2020-WIDESPREAD-2020-5) under grant agreement No. 951737 (NONGAUSS). O.L. and L.M. acknowledge the project 8C22002 (CVStar) of MEYS of the Czech Republic, which has received funding from the European Union’s Horizon 2020 research and innovation programme under the Grant Agreements no. 731473 and 101017733; and the project ’Quantum Secure Networks Partnership’ (QSNP), which has received funding from the European Union’s Horizon Europe research and innovation programme under the Grant Agreement no. 101114043.
@@ Appendix A Fully inseparable biseparable Gaussian state
This section is dedicated to the proof that the three-mode biseparable state (3) with CM (12) is in a certain region of squeezing parameters fully inseparable, i.e., it is entangled across all three bipartite splits $1|23,2|31$ and $3|12$. The state possesses a fully symmetric CM of the form
@@
of modes $1$ and $3$. Entanglement of the state with the latter CM then can be proved easily by using positive partial transposition criterion [54] expressed in terms of the symplectic eigenvalues [74]. The partial transposition operation with respect to mode $3$ transforms the CM (85) as $\gamma_{13}^{{\rm test}\,(T_{3})}=(\mathbbm{1}\oplus\sigma_{z})\gamma_{13}^{\rm test}(\mathbbm{1}\oplus\sigma_{z})$ and its symplectic eigenvalues $\mu_{\pm}$ are obtained from the spectrum $\{\pm i\mu_{+},\pm i\mu_{-}\}$ of the matrix $\Omega_{2}\gamma_{13}^{{\rm test}\,(T_{3})}$ [74] in the form
@@
Appendix B Derivation of the structure of the matrix $Z_{A}^{\alpha}$ providing two-mode quadrature combinations $u_{i}^{\alpha}$
In this section we show what structure the matrix $Z_{A}^{\alpha}$ must have for the operators $u_{i}^{x,p}$, $i=1,2,\ldots,N-1$, Eq. (19) of the main text, to contain at most two non-zero terms. In the main text, we showed that for operators $u_{i}^{x,p}$ to contain at most two non-zero terms, the vectors $L^{\alpha}_{i+1:N,i}$, $i=1,2,\ldots,N-1$, Eq. (23) of the main text, have to possess at most one non-zero component. Below we show that the vectors $L^{\alpha}_{i+1:N,i}$, will have at most one non-zero component if the vectors $(Z_{A}^{\alpha})_{i+1:N,i}$, $i=1,2,\ldots,N-1$, contain at most one non-zero component.
Namely, for a positive-definite matrix $Z_{A}^{\alpha}$ the vectors $L^{\alpha}_{i+1:N,i}$ can be expressed as [63, Theorem 4.2.5]
$L^{\alpha}_{2:N,1}=\frac{(Z_{A}^{\alpha})_{2:N,1}}{L_{11}^{\alpha}}$ (88)
and
$L^{\alpha}_{i+1:N,i}=\frac{1}{L_{ii}^{\alpha}}\left[(Z_{A}^{\alpha})_{i+1:N,i}-\sum_{j=1}^{i-1}L_{ij}^{\alpha}L^{\alpha}_{i+1:N,j}\right]$ (89)
for $i=2,3,\ldots,N-1$. From Eq. (88) it then follows that the vector $L^{\alpha}_{2:N,1}$ will contain at most one non-zero component if the vector $(Z_{A}^{\alpha})_{2:N,1}$ has at most one non-zero component. Moving to the vector $L^{\alpha}_{3:N,2}$, Eq. (89) with $i=2$, we see that if the vector $(Z_{A}^{\alpha})_{2:N,1}$ has at most one non-zero component, then the second term in the square brackets vanishes because it contains products of different components of the vector $(Z_{A}^{\alpha})_{2:N,1}$, and $L^{\alpha}_{3:N,2}=(Z_{A}^{\alpha})_{3:N,2}/L^{\alpha}_{22}$ according to Eq. (89). Again, in order the vector $L^{\alpha}_{3:N,2}$ to have at most one non-zero component, the vector $(Z_{A}^{\alpha})_{3:N,2}$ has to possess at most one non-zero component. Repeating the same procedure (i1)(i-1)-times we see that the vectors $L^{\alpha}_{j+1:N,j}$, $j=1,2,\ldots,i-1$ will have at most one non-zero component if the same holds for vectors $(Z_{A}^{\alpha})_{j+1:N,j}$, $j=1,2,\ldots,i-1$. If this is the case, then the sum $\sum_{j=1}^{i-1}L_{ij}^{\alpha}L^{\alpha}_{i+1:N,j}$ in Eq. (89) vanishes, because each term $L_{ij}^{\alpha}L^{\alpha}_{i+1:N,j}$ is a vector with components given by a product of different components of the vector $(Z_{A}^{\alpha})_{j+1:N,j}$. The $i$-th vector (89) of the Cholesky matrix $L^{\alpha}$ then reduces to
$L^{\alpha}_{i+1:N,i}=\frac{(Z_{A}^{\alpha})_{i+1:N,i}}{L_{ii}^{\alpha}}$ (90)
and again it has at most one non-zero component if the vector $(Z_{A}^{\alpha})_{i+1:N,i}$ contains at most one non-zero component.
Summarizing the obtained results we see that if we require all the quadrature combinations $u_{i}^{x,p}$, Eq. (19) of the main text, to contain at most two non-zero terms, then all vectors $(Z_{A}^{\alpha})_{i+1:N,i}$, $i=1,2,\ldots,N-1$ have to possess at most one non-zero component, as we wanted to prove.
Appendix C Optimal parameters
@@
For the three-mode CM (64) and the product criterion (5.1) the optimal parameters $\ell^{\alpha}_{ij}$ are summarized in Tab. 4
Table 4: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference $D_{P}=-0.245$ for the product criterion (5.1) and the CM $\gamma_{1}$, Eq. (64).
@@
In the case of CM (64) and the sum criterion (5.1) the optimal parameters $\ell^{\alpha}_{ij}$ are summarized in Tab. 5
Table 5: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference $D_{S}=-0.165$ for the sum criterion (5.1) and the CM $\gamma_{1}$, Eq. (64).
@@
Optimal parameters $\ell^{\alpha}_{ij}$ for the CM (65) and the sum criterion (5.1) are summarized in Tab. 6
Table 6: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference $D_{S}=-0.070$ for the sum criterion (5.1) and the CM $\gamma_{2}$, Eq. (65).
@@
$\ell^{x}_{32}$ -0.908 $\ell^{p}_{32}$ 0.559
[…]
Appendix D Derivation of examples from a guessed witness